OpenStudy (anonymous):

kFind an equation for the line perpendicular to 4x+40y=360and goes through the point (-1,9 ). Write your answer in the form y=mx+b .

OpenStudy (anonymous):

divide through the equation by 40. we want to isolate or solve for y because we need to put it in slope intercept form (y=mx+b)

OpenStudy (anonymous):

i did and i got y=4/40x+9 and i reduce it and got y=1/10x+9

OpenStudy (anonymous):

i think you may have overlooked a negative sign. can you see that?

OpenStudy (anonymous):

yes. so it will be y=-1/10x-9

OpenStudy (anonymous):

close. the 9 stays pos.

OpenStudy (anonymous):

so you get y=-1/10 x+9.

OpenStudy (anonymous):

then i use that slope to find the equation for this points (-1,9)

OpenStudy (anonymous):

yes. so what is the slope and what is the slope of the perpendicular line?

OpenStudy (anonymous):

so it would be y=-1/10x+8

OpenStudy (anonymous):

the first line has the equation y=-1/10x+9

OpenStudy (anonymous):

now what do you think the slope of that line is?

OpenStudy (anonymous):

positive 1/10

OpenStudy (anonymous):

because from that slope, we us the rule that two lines that are perpendicular have slopes that are negative reciprocals of eachother.

OpenStudy (anonymous):

you are close.

OpenStudy (anonymous):

the slope is actually -1/10

OpenStudy (anonymous):

y=-1/10x+8

OpenStudy (anonymous):

whatever the number is in front of the x, that is the slope in slope intercept form

OpenStudy (anonymous):

we know that the slope of the original line is-1/10, so what is the slope of the perpendicular line?

OpenStudy (anonymous):

10/1

OpenStudy (anonymous):

nice. absolutely correct.

OpenStudy (anonymous):

so now, you can substitute the coordinates (-1,9) into the equation y=mx+b.

OpenStudy (anonymous):

you should e getting something like 9=10(-1)+b.

OpenStudy (anonymous):

solving for b gives b=20. so your equation is y=10x+20

OpenStudy (anonymous):

so it would be y=10x+1

OpenStudy (anonymous):

in order to solve 9=10(-1)+b, you get 9=-10+b; then add 10 to both sides of the equation to get b=20. so your equation is y=10x+20

OpenStudy (anonymous):

feel free to ask a question if something is unclear.

OpenStudy (anonymous):

i get it. thanks

OpenStudy (anonymous):

no problem. take care

OpenStudy (anonymous):

you to